OpenStudy (anonymous):

One CD player is 3 times as much as another. Together the two CD players would be $320. What is the price of the more expensive CD player?

OpenStudy (anonymous):

a = 3*b a + b = 320 You substitute and get: 3b + b =320, so b = 80, and therefore a = 320 - 60 = 180 The most expensive one costs 180$

OpenStudy (anonymous):

240$

OpenStudy (anonymous):

Oops ^^ sorry, I mistyped, i meant a = 320 - 80 = 240